# A 2-step Framework for Automated Literary Translation Evaluation: Its Promises and Pitfalls

### v2

@@ F.1 Baseline Details
Appendix G Prompting Methods Used for Generating Diverse Translations
To ensure that we have a diverse set of translations for the human evaluation stage mentioned in Section 5, we took several big and small language models and prompted them in various ways to obtain translations. We specifically used a combination following prompting techniques:
  1. 1.
    Sentence-wise prompting where we tokenize each paragraph into individual sentences and then give it to LLMs to translate.
  2. 2.
    Paragraph level prompting where we give LLMs the whole paragraph and ask it to translate in the target language.
  3. 3.
    We also incorporate zero shot and few shot prompting for the abovementioned scenario. Human inspection revealed that providing a few example translations always leads to better translation, as echoed in previous work like Zhang et al. (2023).
  4. 4.
    In a few variants we also add an LLM generated summary on top of few shot examples.
For the final analysis in Section 6 we generate translations using five-shot prompting with GPT4o summary appended. The five examples were sourced from stories other than the one that is being translated. We made sure to keep both dialogue-based and narrative-based paragraphs in the few shot examples.
All LLM translations were generated using the default temperature of the model (as opposed to 0.0 for evaluation).
Appendix H Additional Results and Ablations
H.1 LLM Generated Questions to Evaluate Literary Translations
H.1.1 Categories
@@
"Historical and Cultural Context", "Imagery and Descriptive Quality", "Character Voice, Tone, and Individuality", "Interpersonal Communication and Hierarchy", "Linguistic and Idiomatic Naturalness", "Nuanced Interpretation including Subtle Implications", "Narrative Pacing and Rhythm", "Affective and Stylistic Resonance", "Overall Consistency and Cohesion"
@@
H.1.2 Quality of the LLM Generated Questions
On average, GPT4-o generates 9.6$\pm$3.4 questions per paragraph. A careful examination by a human expert on 50 randomly selected questions finds 44 questions to be “very helpful" to evaluate literary quality. The six remaining questions, while still relevant, are deemed either “too generic" (4) or unanswerable based on the context of that paragraph alone (2). For example, GPT4-o generates multiple questions about the historical and cultural context for a story set in the mid-19th century America ("The Passing of Grandison"), whereas it generates questions about nuanced interpretations of certain phrases or the use of specific imagery for a story with rich poetic descriptions ("The Ones Who Walk Away from Omelas").
H.1.3 Zero Shot Classification Accuracy
We use GPT4-o to classify the questions. We passed all the questions belonging to a particular paragraph to the LLM and got them classified together to save token cost. Note that, in a few cases ( 5%) the pipeline failed to generate questions that fit the formatting requirement for the classification pipeline and we discarded those questions from our analysis. In most of these cases, the question generator model generated something that couldn’t be parsed properly. This is, however, not a fundamentally difficult issue to solve.
Figure 5 shows the confusion matrix of GPT4o’s ability to classify questions. Here the average accuracy was 0.77 with precision = 0.76, recall = 0.78 and f1 = 0.75.
H.2 Detailed Analysis of each criterion of Step 1
We observe some interesting patterns across the score distribution for Step 1 RULER, especially for Korean honorifics.
@@ H.2 Detailed Analysis of each criterion of Step 1
Figure 6: The confusion matrices show the scores given by two human annotators and LLM for Korean Honorific. As the upper row shows there is good agreement between the annotators that some paragraphs have very poor quality in terms of honorifics (notice the dark blue corner on the upper left). However, as the second row shows, GPT4-o fails to capture this (the upper left corner is lighter). The matrices are normalized column-wise (axis=0)
The Case of Korean Honorific
Honorific is one of the most important aspects of Korean translation, however, honorifics are only relevant in dialogue-like texts. If the text does not involve dialogue, the honorific score is meant to be given 5 according to our rubric. This means that there were a large number of evaluation data points where honorific was not important.
In Korean, the reading experience is heavily affected if proper honorific is not used, it is crucial for the translation judge to be able to capture those mistakes. However, the best models’ precision and recall for low honorific labels is very low (see the bottom row of Figure 6). Table 3 shows aggregated F1 values for all the criteria for each label and GPT4-o has poor F1 values for scores 1, 2, and 3.
Figure 7: Each row represents an agreement between annotator 1 and 2 and comparison with annotator 2’s agreement (left) with GPT4o for each criterion (right). Notice that GPT4-o is better at matching the diagonal patterns in human agreement for these criteria.
Notice that in Figure 7, the language model is better at matching the preference distribution of humans (which is almost diagonal) in other dimensions.
H.3 Details for of VERSE (Step 2)
Table 8 shows classification statistics averaged over three annotators for each score class. Models tend to do worse when humans gave a Score 1, that is, models struggle to spot a bad translation. This is similar to the behavior observed for honorifics in RULER (Step 1).
Human-Human Human-Model
Label Support Prec Rec F1 Prec Rec F1
1 197/189 0.81 0.75 0.78 0.69 0.06 0.11
2 182/171 0.52 0.46 0.49 0.32 0.78 0.45
3 185/204 0.64 0.77 0.70 0.68 0.45 0.54
Avg 188/188 0.66 0.66 0.66 0.56 0.43 0.42
Table 8: Comparison of classification metrics between human annotators (Human-Human) and GPT4-o no CoT predictions (Human-Model). Support shows the average number of instances for human/model evaluations.
H.4 Ablations for RULER (Step 1)
CoT hurts across all models
Table 9 of Appendix H, demonstrates chain of thought prompting seems to hurt performance across all models, except for some models which see a minor improvement in honorifics when used CoT. This observation is at odds with the common practice of LLM-as-a-judge literature and deserves further study.
Honorifics Lexical Syntax Content Aggregated Avg.
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE Avg. $\tau$ Avg. $\rho$ Avg. MSE
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
w/ CoT .44 .48 2.00 .62 .71 0.86 .59 .68 1.02 .61 .70 1.12 .56 .64 1.25
claude-3-5-sonnet .39 .41 2.30 .58 .65 1.17 .57 .65 1.20 .57 .64 1.38 .53 .59 1.51
w/ CoT .45 .48 2.15 .53 .61 1.21 .55 .63 1.51 .56 .63 1.58 .52 .59 1.61
claude-3-5-haiku .41 .44 1.86 .54 .60 1.37 .57 .64 1.23 .55 .62 1.42 .52 .58 1.47
w/ CoT .44 .47 1.92 .53 .60 1.24 .46 .53 1.62 .32 .36 1.91 .44 .49 1.67
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
w/ CoT .28 .31 3.36 .46 .51 1.51 .47 .54 1.36 .60 .69 1.00 .45 .51 1.81
claude-3-haiku .20 .21 2.24 .45 .50 1.57 .42 .47 1.47 .50 .56 1.55 .39 .43 1.71
w/ CoT .16 .17 2.28 .40 .44 1.74 .44 .49 1.61 .39 .43 1.89 .34 .38 1.88
gemma-2-27b .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
w/ CoT .28 .31 2.11 .37 .42 1.46 .35 .40 1.45 .52 .58 1.34 .38 .43 1.59
gemma-2-9b .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
w/ CoT .32 .35 2.01 .44 .49 1.35 .29 .33 1.65 .31 .35 1.59 .34 .38 1.65
llama-3-8b -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
w/ CoT .12 .13 2.50 .13 .14 1.90 .20 .23 1.82 .26 .29 1.84 .18 .20 2.01
Table 9: Performance comparison across different linguistic aspects with and without Chain-of-Thought (CoT) prompting. Gray rows indicate results with CoT prompting. $\tau$: Kendall’s Tau, $\rho$: Spearman’s correlation, MSE: Mean Squared Error.
Specific Rubric is Important
As shown in Table 10 providing fine-grained human-curated rubric is important for achieving high correlation with GPT4-o. If the detailed rubric is removed. However, different models are affected differently and for some models having a rubric is critical.
Honorifics Lexical Syntax Content Aggregated Avg.
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE Avg. $\tau$ Avg. $\rho$ Avg. MSE
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
w/o rubric .22 .26 4.67 .68 .76 0.81 .62 .71 0.88 .65 .73 0.85 .54 .61 1.80
w/o human TR .50 .53 2.04 .68 .77 0.82 .65 .75 1.10 .69 .77 1.01 .63 .71 1.24
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
w/o rubric .20 .22 3.36 .63 .70 0.92 .62 .69 0.93 .62 .69 0.96 .52 .58 1.54
w/o human TR .30 .34 2.64 .61 .68 0.94 .58 .66 0.99 .60 .68 1.17 .52 .59 1.43
gemma-2-27b-it .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
w/o rubric .31 .33 3.96 .51 .57 1.50 .53 .60 1.17 .58 .66 1.10 .48 .54 1.93
w/o human TR .37 .40 1.96 .52 .58 1.12 .52 .59 1.09 .54 .62 1.17 .49 .55 1.34
gemma-2-9b-it .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
w/o rubric .20 .21 3.87 .27 .30 1.81 .26 .29 1.72 .20 .22 1.77 .23 .25 2.29
w/o human TR .29 .32 2.15 .54 .60 1.37 .49 .54 1.53 .51 .57 1.54 .46 .51 1.65
llama-3-8b-it -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
w/o rubric -.04 -.05 3.21 .17 .19 2.03 .13 .15 2.19 .15 .17 2.23 .10 .11 2.42
w/o human TR .11 .12 2.39 .36 .40 1.65 .31 .35 1.74 .15 .17 2.23 .23 .26 2.00
@@ H.4 Ablations for RULER (Step 1)
Table 10: (Step 1) Ablation study results comparing model performance under different conditions. Base model performance is shown in white rows, while ablation conditions (without rubric and without human training) are shown in gray rows. Note that for most models, removing the rubric has a bigger effect than removing the human translation
Human Translation is Not Necessary But Good to Have
As seen in Table 10, while human rubrics are good to have, removing them doesn’t lead to a significant performance drop (in terms of correlation) in most criteria. In fact, some models do slightly better without human translation. This is good news since having a human translation for literary text is not possible for most of cases. This might also open the door to evaluating different versions of human translations and quantitatively evaluate their differences. However, note that Avg. MSE is higher for w/o human reference cases, which usually means LLMs are biased towards giving higher scores to machine translation when human translation is not provided.
Honorifics Lexical Syntax Content Aggr. Avg.
Model $k$ $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE Avg. $\tau$ Avg. $\rho$ Avg. MSE
gpt-4o 0 .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
5 .49 .52 1.98 .68 .77 0.75 .63 .72 0.96 .67 .76 0.79 .62 .69 1.12
10 .49 .53 2.00 .69 .78 0.75 .64 .73 0.97 .65 .73 0.88 .62 .69 1.15
15 .48 .52 2.15 .68 .78 0.75 .64 .74 1.02 .66 .75 0.85 .62 .70 1.19
20 .44 .48 2.48 .67 .76 0.80 .64 .73 0.95 .66 .75 0.89 .60 .68 1.28
gpt-4o-mini 0 .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
5 .30 .33 3.40 .54 .61 1.19 .60 .67 0.93 .59 .66 1.02 .51 .57 1.64
10 .30 .33 3.31 .52 .59 1.31 .53 .60 1.10 .59 .66 1.01 .49 .55 1.68
15 .33 .37 3.14 .48 .54 1.36 .50 .57 1.17 .64 .71 0.92 .49 .55 1.65
20 .25 .27 3.16 .51 .57 1.36 .51 .58 1.14 .59 .66 1.01 .46 .52 1.67
Llama-3.1-8B-It 0 -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
5 .19 .20 2.32 .20 .23 1.97 .12 .14 2.21 .23 .26 2.03 .19 .21 2.13
10 .10 .10 2.25 .13 .14 2.05 .14 .16 2.22 .14 .16 2.17 .13 .14 2.17
15 .03 .03 2.32 .18 .20 2.01 .16 .18 2.19 .14 .16 2.19 .13 .14 2.18
20 .24 .26 2.22 .24 .27 1.89 .19 .22 2.08 .23 .26 2.04 .23 .25 2.06
Table 11: (Step 1) Ablation showing the effect of varying $k$ (number of demonstrations) on model performance across different metrics (no chain of thought). Notice that for large models, providing examples doesn’t make a significant difference. For the small Llama-3.1 model, the pattern is more inconsistent, but no significant gain was observed either. $\tau$: Kendall’s Tau, $\rho$: Spearman’s correlation, MSE: Mean Squared Error. Note that gemma models were excluded from this analysis because of their shorter context windows. Human translations were evaluated in a no-reference format.
Providing few shot examples doesn’t help
We experimented with few shot prompting by providing a few examples (without CoT). However, as the results in Table 11 show there is no significant performance gain across models when few annotated examples were given. In fact, some models tend to do better without few shot examples. We kept the few shot examples the same across all models and made sure that there was at least one example from each score (1 to 5) among the examples.
H.5 Ablations for VERSE (Step 2)
Model $\tau$ $\rho$ MSE
gpt-4o 0.49 0.54 0.61
w/ CoT 0.45 0.50 0.69
claude-3-5-sonnet 0.51 0.54 0.73
w/ CoT 0.36 0.39 0.65
claude-3-5-haiku 0.37 0.40 0.77
w/ CoT 0.39 0.42 0.67
gpt-4o-mini 0.37 0.40 0.61
w/ CoT 0.34 0.36 0.63
claude-3-haiku 0.20 0.21 0.82
w/ CoT 0.17 0.19 0.89
gemma-2-27b 0.13 0.14 0.71
w/ CoT 0.09 0.10 0.78
gemma-2-9b 0.09 0.09 0.71
w/ CoT 0.16 0.17 0.88
llama-3-8b 0.05 0.05 1.08
w/ CoT 0.14 0.16 0.90
Table 12: (Step 2) Performance of different models with and without Chain-of-Thought (CoT) prompting. The effect of CoT is inconsistent and depends on the specific models. Some models do benefit from CoT.
CoT only helps some models
As seen in Table 12, chain of thought reasoning doesn’t GPT series models much, in fact, it hurts. However, for smaller models, doing chain of thought boosts the performance from completely random to 0.20.2~{}0.20.2. However, even for the models that benefit from CoT (in terms of correlation), some of them have higher MSE, which usually means they tend to give higher scores to machine translations.
Default No Rubric No Human Tr
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE
gpt-4o 0.49 0.54 0.61 0.08 0.09 1.22 0.52 0.57 0.59
gpt-4o-mini 0.37 0.40 0.61 0.02 0.02 1.05 0.42 0.46 0.60
gemma-2-27b 0.13 0.14 0.71 0.07 0.08 0.70 0.25 0.26 0.75
gemma-2-9b 0.09 0.09 0.71 0.02 0.02 0.79 0.17 0.18 0.75
llama-3-8b 0.05 0.05 1.08 -0.02 -0.02 0.73 0.06 0.06 1.08
Table 13: (Step 2) Performance comparison across different models and configurations. No Rubric refers to when human written rubric was not provided and No Human Tr means human reference translation was not provided. Notice that while having access to human translation is not critical for large models, having access to human-curated rubic is essential for all models.
Extensive rubric is critical but reference translation is not
As seen in Table 13, not providing human written criteria leads the correlation to drop to zero. A closer inspection revealed that when a specific grading rubric is not provided, the model gives most of the translations a 3 out of 3. While the number of 2s doesn’t change much, not providing specific rubrics decreases the number of 1s the model provides.
The table also shows that not providing human translation works better than providing reference translation.
Model k $\tau$ $\rho$ MSE
gpt-4o 0 0.49 0.54 0.61
5 0.47 0.51 0.65
10 0.52 0.57 0.58
15 0.48 0.53 0.61
20 0.50 0.54 0.63
gpt-4o-mini 0 0.37 0.40 0.61
5 0.20 0.21 0.75
10 0.14 0.15 0.74
15 0.20 0.22 0.73
20 0.19 0.21 0.77
llama-3-8b-it 0 0.05 0.05 1.08
5 0.04 0.04 1.59
10 0.06 0.07 1.31
15 0.06 0.06 0.82
20 0.05 0.05 0.77
Table 14: (Step 2) Performance comparison across k shot examples provided (in no COT setting) in answering LLM generated evaluation questions. Note that gemma models are excluded because of their shorter context length.
Providing few shot examples doesn’t make a big difference
As Table 14 shows, providing a few shot examples doesn’t significantly boost performance across models.
Appendix I Additional Results from Discussion (Section 6)
I.1 Translation Generation
We selected the bottom 10 stories from Table 7 and generated their translations using language models and Google Translate. We use the top 5 stories of Table 7 to obtain few shot examples. We used one paragraph from each story for few shot prompting. One example paragraph translated across different models can be found in Table 17.

### v3

@@ B.1 Baseline Details
GEMBA-MQM:
We used the reference free mode of GEMBA-MQM Kocmi and Federmann (2023) since the original paper claims it to be superior to the mode using reference translation. We used the same GPT4o model that powers most of our experiments.
Note that GEMBA-MQM and our framework’s numbers are not really comparable since the MQM categories and our categories are very different.
Appendix C Prompting Methods Used for Generating Diverse Translations
To obtain diverse translations for the human evaluation stage (Section 5), we explored several prompting setups with both large and small LLMs. Translations were generated either sentence by sentence or at the paragraph level, and we varied between zero-shot and few-shot settings. Consistent with prior work Zhang et al. (2023), as well as our own inspection, few-shot prompting generally produced higher-quality outputs. In some cases we also prepended an LLM-generated summary to the few-shot examples. For the final analysis (§6), we used five-shot prompting with a GPT-4o summary. The examples were drawn from stories other than the one being evaluated. We made sure to include both dialogue- and narrative-based paragraphs. All translations were generated with the models’ default temperature (unlike evaluation, which used $T=0.0$).
C.1 LLM Generated Questions to Evaluate Literary Translations
C.1.1 Categories
@@
”Historical and Cultural Context”, ”Imagery and Descriptive Quality”, ”Character Voice, Tone, and Individuality”, ”Interpersonal Communication and Hierarchy”, ”Linguistic and Idiomatic Naturalness”, ”Nuanced Interpretation including Subtle Implications”, ”Narrative Pacing and Rhythm”, ”Affective and Stylistic Resonance”, ”Overall Consistency and Cohesion”
@@
Appendix D Details of Common Issues found in Machine Translation
Word/Phrase/Expression Level Issues
This type of error occurs most frequently and significantly impacts overall translation quality. While LLMs excel at literal translations, they struggle with natural word choices in the target language, particularly when dealing with figurative expressions. This often results in the selection of awkward or uncommon verbs in Korean; although the overall content remains comprehensible, these choices critically diminish the quality of translation. Notably, problems often arise with polysemous words or idiomatic expressions, seriously impeding reader comprehension. Additional common errors include inappropriate use of pronouns, such as referring to people as “그것 (it),” reflecting the essential difference in the linguistic structure between English and Korean.
Context Related Issues
This type of error presents more complex and serious challenges arising from cultural and historical differences between the source and target languages. For instance, because terms like “abolitionist” in “The Passing of Grandison” do not have an equivalent historical term in Korean language, LLMs end up offering incomprehensible expressions. Similarly, the expression “border states” often loses its geographical context (mistranslated as “국경주–states close to national borders” in Korean). LLMs also struggle with specialized terminology and historical references, as shown in mistranslations of terms like “bootjack” and “nightcap.”
Form-related issues
This third category includes subcategories concerning punctuation and grammar, as well as structure and formatting. Punctuation and grammar refer to problems with punctuation changes that alter meaning. For example, when LLM translation replaces semicolons with periods, it alters the original text’s meaning. Likewise, changes in the sentence structure in the translated text can lead to incomplete or awkward results. This subcategory includes problems with splitting long sentences unnecessarily, which results in awkward structures when handling complex sentences. Other issues in this category included structure and formatting. These appeared when formatting the dialogue and narrative flows. The incorrect handling of sentence structures in the LLM translation led to awkward sentences, negatively affecting readability. Form-related issues also include transliteration, consistency issues, and omissions. LLMs transliterate Western names inconsistently in the translation, resulting in different name variations throughout the translated text. (i.e., translating “Grandison” as “그라디슨” and “그란디손” within a single translation) LLM translations may eliminate a sentence and expression or add arbitrary content different from the original text. They also leave certain words untranslated, sometimes including various words or phrases in random languages within the translation.
Honorifics
Appendix E Additional Results and Ablations
@@
Figure 5: Confusion matrices for Korean honorifics show strong agreement between human annotators (top row), who consistently mark some paragraphs as very poor (dark blue upper-left corner). In contrast, GPT-4o fails to capture this pattern (lighter upper-left corner, bottom row).
Honorifics Lexical Syntax Content Step 2 VERSE
Model $k$ $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE
gpt-4o 0 .51 .56 2.03 .68 .77 0.72 .62 .71 0.97 .70 .78 0.81 0.49 0.53 0.63
5 .49 .53 1.91 .68 .77 0.76 .62 .71 0.98 .68 .76 0.79 0.47 0.52 0.64
10 .50 .53 1.93 .70 .78 0.73 .64 .73 0.99 .65 .73 0.88 0.51 0.55 0.59
15 .49 .53 2.07 .69 .78 0.73 .64 .73 1.03 .66 .75 0.86 0.48 0.53 0.61
20 .44 .48 2.39 .67 .77 0.80 .64 .73 0.97 .66 .75 0.89 0.49 0.53 0.63
gpt-4o-mini 0 .27 .31 2.74 .63 .70 0.93 .59 .67 0.93 .60 .67 1.08 0.37 0.40 0.63
5 .31 .35 3.28 .55 .62 1.19 .60 .67 0.93 .60 .67 1.00 0.21 0.23 0.75
10 .31 .34 3.20 .54 .61 1.30 .54 .61 1.10 .60 .67 0.99 0.15 0.16 0.75
15 .34 .38 3.03 .49 .56 1.36 .49 .56 1.22 .64 .71 0.90 0.23 0.25 0.72
20 .26 .29 3.05 .53 .59 1.35 .50 .57 1.19 .59 .66 1.01 0.20 0.22 0.78
llama-3-8b-it 0 -.01 -.01 2.31 .37 .41 1.56 .29 .32 1.65 .21 .24 2.04 0.05 0.06 1.06
5 .19 .20 2.27 .21 .23 1.93 .12 .14 2.16 .24 .26 1.98 0.05 0.05 1.53
10 .11 .12 2.19 .13 .14 2.01 .14 .16 2.17 .14 .16 2.12 0.09 0.10 1.27
15 .03 .03 2.27 .18 .20 1.97 .16 .18 2.14 .14 .16 2.14 0.06 0.07 0.82
20 .25 .26 2.18 .25 .28 1.85 .20 .22 2.03 .24 .26 2.00 0.06 0.06 0.77
Table 10: (Performance comparison when k shot examples were provided (in no COT setting) in RULER metrics and VERSE (Step 2). The best results for each model are shown in bold. In most cases providing few shot examples do not consistently boost model performance and for a few cases it even seems to decrease the performance.
Honorifics Lexical Syntax Content Step 2 (VERSE)
Eval Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE
gpt-4o .51 .56 2.03 .68 .77 0.72 .62 .71 0.97 .70 .78 0.81 0.49 0.53 0.63
[…]
gpt-4.1-2025-04-14 .52 .56 1.61 .66 .75 0.78 .64 .73 0.84 .66 .76 0.84 .42 .46 0.66
[…]
claude-3-5-sonnet .40 .41 2.22 .60 .67 1.13 .59 .66 1.17 .57 .64 1.35 0.38 0.41 0.63
[…]
claude-3-5-haiku .42 .45 1.80 .54 .59 1.35 .56 .64 1.22 .55 .62 1.39 0.38 0.41 0.75
[…]
gpt-4o-mini .27 .31 2.74 .63 .70 0.93 .59 .67 0.93 .60 .67 1.08 0.37 0.40 0.63
[…]
claude-3-haiku .22 .23 2.17 .45 .50 1.55 .42 .47 1.47 .51 .56 1.51 0.20 0.21 0.81
[…]
gemma-2-27b .34 .38 2.08 .47 .52 1.27 .49 .56 1.24 .49 .55 1.25 0.18 0.19 0.70
[…]
gemma-2-9b .28 .30 2.36 .43 .48 1.40 .36 .41 1.47 .52 .58 1.34 0.09 0.09 0.73
[…]
llama-3-8b -.01 -.01 2.31 .37 .41 1.56 .29 .32 1.65 .21 .24 2.04 0.05 0.06 1.06
[…]
deepseek-v3 .47 .52 1.75 .66 .74 0.90 .60 .67 0.96 .67 .74 0.95 .51 .55 0.75
[…]
deepseek-r1 .59 .64 1.35 .57 .65 1.05 .57 .66 1.01 .59 .68 1.04 .50 .55 0.65
[…]
o3-2025-04-16 .53 .57 1.40 .65 .73 0.87 .60 .68 0.93 .59 .67 1.31 .43 .47 0.63
[…]
Table 11: Effect of Chain of Thought across Step 1 RULER (Honorifics, Lexical, Syntax, Content) and Step 2 VERSE (Overall Aggregated). Results show that CoT has mixed effects: while most models benefit from CoT in evaluating Honorifics, for most models CoT does not improve performance, and in some cases even hurts. This pattern hold true even for the reasoning models where the models produce long reasoning trace before outputting answer. o3 was evaluated at $T=1.0$, the only available temperature setting.
Honorifics Lexical Syntax Content Step 2 (VERSE)
Eval Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE
gpt-4o .51 .56 2.03 .68 .77 0.72 .62 .71 .97 .70 .78 0.81 0.49 0.53 0.63
[…]
gpt-4o-mini .27 .31 2.74 .63 .70 0.93 .59 .67 0.93 .60 .67 1.08 0.37 0.40 0.63
[…]
gemma-2-27b-it .34 .38 2.08 .47 .52 1.27 .49 .56 1.24 .49 .55 1.25 0.18 0.19 0.70
[…]
gemma-2-9b-it .28 .30 2.36 .43 .48 1.40 .36 .41 1.47 .52 .58 1.34 0.09 0.09 0.73
[…]
llama-3-8b-it -.01 -.01 2.31 .37 .41 1.56 .29 .32 1.65 .21 .24 2.04 0.05 0.06 1.06
[…]
@@
Table 12: Ablation of the effect of human translation and rubric. Note that for most models, removing the rubric has a bigger effect than removing the human translation. In fact, for Step 2 VERSE, the models tend to do better without human translation.